# Kibble-Zurek behavior in a topological phase transition with a quadratic band crossing

### v1

@@ Abstract
Kibble-Zurek (KZ) mechanism describes the scaling behavior when driving a system across a continuous symmetry-breaking transition. Previous studies have shown that the KZ-like scaling behavior also lies in the topological transitions in the Qi-Wu-Zhang model (2D) and the Su-Schrieffer-Heeger model (1D), although symmetry breaking does not exist here. Both models with linear band crossings give that $\nu=1$ and $z=1$. We wonder whether different critical exponents can be acquired in topological transitions beyond linear band crossing. In this work, we look into the KZ behavior in a topological 2D checkerboard lattice with a quadratic band crossing. We investigate from dual perspectives: momentum distribution of the Berry curvature in clean systems for simplicity, and real-space analysis of domain-like local Chern marker configurations in disordered systems, which is a more intuitive analog to conventional KZ description. In equilibrium, we find the correlation length diverges with a power $\nu\simeq 1/2$. Then, by slowly quenching the system across the topological phase transition, we find that the freeze-out time $t_{\mathrm{f}}$ and the unfrozen length scale $\xi(t_{\mathrm{f}})$ both satisfy the KZ scaling, verifying $z\simeq 2$. We subsequently explore KZ behavior in topological phase transitions with other higher-order band crossing and find the relationship between the critical exponents and the order. Our results extend the understanding of the KZ mechanism and non-equilibrium topological phase transitions.
@@ I Introduction
Quantum phase transitions are one of the fascinating topics in modern physics. Conventional continuous phase transitions are driven by spontaneous symmetry breaking and are usually described in terms of local order parameters that take continuous values Miransky (1994); Vojta (2003); Sachdev (2011); Nie and Zheng (2023a). In contrast, topological phase transitions are a new class of transitions beyond Landau’s symmetry breaking theory, characterized by a quantized, non-local topological invariant Hasan and Kane (2010); Qi and Zhang (2011). Recently, the non-equilibrium dynamics of topological phase transitions have received widespread attention both experimentally and theoretically Fläschner et al. (2018); D’Alessio and Rigol (2015); Hauke et al. (2014); Ye and Li (2020); Sun et al. (2018); Nie and Zheng (2023b), particularly concerning the critical behavior of quenching across topological phase transitions Liou and Yang (2018); Ulčakar et al. (2019, 2020); Sun et al. (2022); Kuo et al. (2021).
@@
Relating conventional continuous phase transitions and nonequilibrium dynamics, the Kibble-Zurek (KZ) mechanism was originally proposed by Kibble in his study of the formation of topological defects in the early universeKibble (1976), and it has been applied by Zurek to study phase transitions in condensed matter systems Zurek (1985); Zurek et al. (2005). It elucidates the nonequilibrium properties of a system when it is quenched across a continuous phase transition in a finite time $\tau$, which shows universal scaling. In equilibrium, the theory of critical phenomena states that as the control parameters $V$ are tuned closer to their critical values $V_{\mathrm{c}}$, the correlation length $\xi$ and relaxation time $t_{\mathrm{r}}$ tend to diverge as $t_{\mathrm{r}}\sim|V-V_{\mathrm{c}}|^{-z\nu}$ and $\xi\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $z$ is dynamical exponent and $\nu$ is spatial critical exponent. As the system slowly quenches across the transition, the relaxation time diverges as it approaches the critical point, and the system’s evolution ceases to be adiabatic. The moment when the system no longer evolves adiabatically is referred to as the freeze-out time $t_{\mathrm{f}}$. The freeze-out time and the unfrozen correlation length satisfy $t_{\mathrm{f}}\sim\tau^{z\nu/(1+z\nu)}$ and $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}$. The KZ behavior has been observed in diverse experimental systems, such as cosmic microwave background Bevis et al. (2008), superfluids Chesler et al. (2015); Mathey and Polkovnikov (2010), superconductor Sonner et al. (2015); Garaud and Babaev (2014); Monaco et al. (2006), liquid crystals Chuang et al. (1991), and ultracold atoms Donner et al. (2007); Weiler et al. (2008); Ko et al. (2019); Chen et al. (2011); Lamporesi et al. (2013); Navon et al. (2015); Clark et al. (2016); Keesling et al. (2019); Yi et al. (2020); Huang et al. (2021); Liang et al. (2022).
@@
Although topological transitions can not be described by Landau’s symmetry breaking theory, the KZ-like behavior has already been introduced to topological systems. It is first proposed that effective correlation length can be extracted from Berry curvature profile in momentum space, and it shows power-law divergence at the transition Chen et al. (2017). However, this momentum-space description is not quite intuitive in contrast with the conventional KZ mechanism. Recently, significant breakthroughs have been made in the real-space analysis. It has been found that the local Chern marker (LCM) will show domain-like configurations when the system is disordered Ulčakar et al. (2020). As long as the disorder is weak, the disorder-averaged normalized autocovariance function of LCM is independent of the disorder strength. So the autocovariance function can be regarded as an intrinsic property of the original topological system. Then, the correlation length can be determined in this purely real-space method. During the quench, the domain-like configurations are the analogy to conventional topological defects and the correlation length shows KZ behavior. This method is first proposed in two-dimensional Qi-Wu-Zhang (QWZ) insulators Ulčakar et al. (2020) and then extended to one-dimensional Su-Schrieffer-Heeger (SSH) models Sun et al. (2022) by changing LCM to local winding number. However, since these two models have similar linear dispersion while gap closing, they have yielded the same critical exponents, $\nu=1,z=1$. Therefore, we naturally wonder whether different critical exponents can be obtained.
@@
In this work, we systematically investigate the KZ behavior in the topological transition of a 2D checkerboard model with quadratic band crossing. For every scaling studied, two perspectives are adopted: momentum-space Berry curvature in the clean system, and the real-space local Chern marker in the disordered system. In the former, we mainly focus on the radial distribution of Berry curvature. In the latter, we are concerned with the domain-like configurations of LCM and extract real-space correlation lengths.
@@
Figure 1: (a). Checkboard lattice model. Red (Blue) points denote the A (B) sublattices. NN hopping strength along the black arrows direction (opposite direction) is $J-\mathrm{i}V$ ($J+\mathrm{i}V$), which breaks the time-reversal symmetry as long as $V\neq 0$. Red (Blue) solid lines represent the NNN hopping strength $J_{1}^{\prime}(J_{2}^{\prime})$. (b). Topological phase diagram. Chern number $|C|=1$ acquires an opposite sign to $V$. At the transition, it is a time-reversal invariant semimetal. (c). The band structure of our model is in different regimes. The colors denote the polarization $\langle\sigma_{z}\rangle$. The gap closes with a quadratic dispersion at the transition $V=0$. We can see the band-inversion surface (BIS) (shown as the white curves) is not a circle around symmetric momentum as in the QWZ model Zhang et al. (2018); Sun et al. (2018), which is a fundamental effect induced by the quadratic band crossing.
@@
We find that the length scale, in both perspectives, diverges near the critical point of the topological phase transition in equilibrium, satisfying $\xi\sim|V-V_{\mathrm{c}}|^{-1/2}$. By slowly quenching through the topological phase transition over a quench time $\tau$, we observe that the freeze-out time, defined by the delay time of the relaxation of Hall response in the clean system or the divergence of correlation length in the disordered system, satisfies $t_{\mathrm{f}}\sim\tau^{1/2}$. The corresponding length scale at the freeze-out time $t_{\mathrm{f}}$ satisfies $\xi(t_{\mathrm{f}})\sim\tau^{1/4}$, where the dynamic exponent $z=2$ and spatial critical exponent $\nu=1/2$ differs from the results of linear band crossings. Additionally, we extend the analysis to higher-order dispersion relations beyond quadratic band crossings and find that the KZ power-law relations still hold. The corresponding critical exponents are related to the band-crossing relations.
@@ II Model
Two-dimensional (2D) checkerboard lattice Sun et al. (2009, 2011) is an example of a topological model with quadratic band crossing. We consider a simplified checkerboard lattice model consisting of two sublattices (A and B) with real and isotropic nearest-neighbor (NN) hoppings ($J$), two kinds of next-nearest-neighbor (NNN) hoppings ($J_{1}^{\prime},J_{2}^{\prime}$), and complex NN hopping ($\mathrm{i}V$) leading to a topological nontrivial phase (see Fig.1(a)). The Hamiltonian can be written as:
$\begin{split}\hat{H}&=\sum_{\mathbf{r}}|x,y\rangle\langle x,y|\otimes\begin{pmatrix}0&J-\mathrm{i}V\\ J+\mathrm{i}V&0\end{pmatrix}\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x+1,y|\otimes\begin{pmatrix}J_{2}^{\prime}&0\\ J-\mathrm{i}V&J_{1}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x,y+1|\otimes\begin{pmatrix}J_{1}^{\prime}&J+\mathrm{i}V\\ 0&J_{2}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y+1\rangle\langle x+1,y|\otimes\begin{pmatrix}0&0\\ J+\mathrm{i}V&0\end{pmatrix}+\text{H.c.}\right]\\ \end{split}$ (1)
where $\mathbf{r}=(x,y)$ denote the unit cell position (unit cell is shown as a dashed ellipse in Fig.1(a)), and the $2\times 2$ matrices represent the sublattice hopping corresponding to the unit cell hopping.
After Fourier transformation, we can express the Hamiltonian in momentum space ^$\hat{\mathcal{H}}_{\mathbf{k}}$ with momentum $\mathbf{k}=(k_{x},k_{y})$ as
^$\begin{split}\hat{\mathcal{H}}_{\mathbf{k}}=&\left[-\left(J_{1}^{\prime}+J_{2}^{\prime}\right)\left(\cos k_{x}+\cos k_{y}\right)\right]\hat{I}\\ &-4J\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\right)\hat{\sigma}_{x}-2V\left(\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\right)\hat{\sigma}_{y}\\ &-\left(J_{1}^{\prime}-J_{2}^{\prime}\right)\left(\cos k_{x}-\cos k_{y}\right)\hat{\sigma}_{z}\end{split}$ (2)
where $\hat{I}$ and $\hat{\sigma}_{x,y,z}$ are the identity and Pauli matrices, respectively. We set $J=1$, $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}$ and periodic boundary condition for entire text afterwards. It is easy to read from Eq.2 that the system is an insulator without time-reversal symmetry as long as $V\neq 0$. Actually, the system has two topologically non-trivial phases, whose Chern number $C$ is determined by the sign of complex hopping $V$, as shown in Fig.1(b). Additionally, the energy band structures in different regimes are depicted in Fig.1(c). We can see that the energy gap closes at $\mathbf{G}=(k_{x},k_{y})=(\pi,\pi)$ when $V=V_{\mathrm{c}}=0$, exhibiting a quadratic band crossing. This is totally different from the QWZ model. There, the gap closes at a Dirac cone, and every Dirac cone changes the Chern number with $1$. Besides, the band-inversion surface (defined as $\langle\sigma_{z}\rangle=0$ contour) in the checkerboard lattice model forms a crisscross, while in the QWZ model, it is a circle around the band crossing point.
@@ III Ground state
For conventional second-order phase transitions driven by spontaneous symmetry breaking, the correlation lengths exhibit a universal scaling behavior as the systems approach the critical point. When it comes to topological transitions, such as the QWZ model and the SSH model, a scaling of $\nu=1$ has been found. However, can this exponent be different in other models with different band crossings?
@@ III.1 Clean system
Figure 2: (a). Correlation length in the ground state could be determined by the FWHM of the radial distribution of Berry curvature with $J^{\prime}=1$. The closer to the transition, the width of the peak decreases. (b). This correlation length diverges in a power $\nu\simeq 1/2$ as the system approaches the phase transition.
@@
Some studies have provided correlation functions for topological phases, defined as the Fourier transform of the Berry curvature Chen et al. (2017); Chen and Schnyder (2019), which measures the itinerant-circulation correlation between Wannier states. Thonhauser et al. (2005); Marzari et al. (2012). More directly, an alternative definition of correlation length can also be extracted from the profile of the Berry curvature in momentum space. Similar procedures have been used in studying the critical behavior in the QWZ insulators and the SSH model Ulčakar et al. (2020); Sun et al. (2022).
@@
The occupied-band Berry curvature $F(\mathbf{k})$ can be expressed via Berry connection $A_{k_{\alpha}}=\left\langle u_{\mathbf{k}}\left|\mathrm{i}\partial_{k_{\alpha}}\right|u_{\mathbf{k}}\right\rangle$ ($\alpha=x,y$) as
$F(\mathbf{k})=\partial_{k_{x}}A_{k_{y}}-\partial_{k_{y}}A_{k_{x}}$ (3)
@@
where $\left|u_{\mathbf{k}}\right\rangle$ is the occupied Bloch wavefunction at momentum $\mathbf{k}$. The Chern number mentioned before can be calculated by $C=\frac{1}{2\pi}\int_{\rm{FBZ}}F(\mathbf{k})\rm{d}^{2}\mathbf{k}$, where ”FBZ” represents the first Brillouin zone.
@@
Then, by integrating the Berry curvature angularly, $f(k)=\frac{1}{k}\int_{|\mathbf{k}^{\prime}-\mathbf{G}|=k}F(\mathbf{k}^{\prime})\text{d}k^{\prime}$, we obtain its radial distribution $f(k)$. Fig 2(a) illustrates fixed $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=1.5$ with six different values of complex NN hoppings strength $V$. It is evident that the radial distribution of the Berry curvature exhibits a peak. Initially, this peak width increases with the increasing $V$, reaching its maximum at the topological phase transition point $V=0$, and subsequently increases. We can find that there may exist critical behavior near the topological phase transition point.
@@
To describe this critical behavior quantitatively, we define the correlation length $\xi_{k}$ by the inverse of the full width at half maximum (FWHM) value of the radial distribution of Berry curvature $f(k)$. A similar approach has been utilized in the QWZ insulators and the SSH model Ulčakar et al. (2020); Sun et al. (2022). As shown in Fig.2(b), we observe that as the system approaches the critical point $V=0$, the correlation length diverges. We plot the relationship between the correlation length and $V-V_{\mathrm{c}}$ in log-log coordinates under three sets of parameters, namely $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=0.5$, $1$, $1.5$. We find that the correlation length satisfies $\xi_{k}\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/2$. This critical exponent value is distinct from the case of linear dispersion, such as in the QWZ insulators where $\nu\simeq 1$.
@@
Figure 3: (a). The real-space distribution of local Chern marker in a disordered sample with size $N\times N=50\times 50$. The area of the fragment domain increases to the order of the whole system sizes at the transition. (b). Divergence of real-space correlation length $\xi_{r}$ at transition with different system sizes in a log-log scale. Each data is an average of $20$ disorder samples. For a certain system size $N$, the correlation length shows good power law in $64<\xi<N/$ 6 regime (emphasized by the markers). The red line represent the $\left|V-V_{\mathrm{c}}\right|^{-1/2}$ scaling for ref. (c). The fitted power $\nu$ approaches $1/2$ as the system size increases. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
III.2 Disordered system

### v2

@@ Abstract
Kibble-Zurek (KZ) mechanism describes the scaling behavior when driving a system across a continuous symmetry-breaking transition. Previous studies have shown that the KZ-like scaling behavior also lies in the topological transitions in the Qi-Wu-Zhang model (2D) and the Su-Schrieffer-Heeger model (1D), although symmetry breaking does not exist there. Both models with linear band crossings give that $\nu\simeq 1$ and $z\simeq 1$. A natural question is whether different critical exponents can emerge in topological transitions beyond the linear band crossing. In this work, we look into the KZ behavior in a topological 2D checkerboard lattice with a quadratic band crossing. We investigate from dual perspectives: momentum distribution of the Berry curvature in clean systems for simplicity, and real-space analysis of domain-like local Chern marker configurations in disordered systems, which is a more intuitive analog to conventional KZ description. In equilibrium, we find the correlation length diverges at the transition with a power $\nu\simeq 1/2$. Then, by slowly quenching the system across the topological phase transition, we find that the freeze-out time $t_{\mathrm{f}}$ and the unfrozen length scale $\xi(t_{\mathrm{f}})$ both satisfy the KZ scaling, verifying $z\simeq 2$. We subsequently explore KZ behavior in topological phase transitions with other higher-order band crossings and find the relationship between the critical exponents and the order. Our results extend the understanding of the KZ mechanism and non-equilibrium topological phase transitions.
@@ I Introduction
Quantum phase transitions are one of the fascinating topics in modern physics. Conventional continuous phase transitions are driven by spontaneous symmetry breaking and are usually described in terms of local order parameters that take continuous values Miransky (1994); Vojta (2003); Sachdev (2011); Nie and Zheng (2023a); Jara and Cosme (2024). In contrast, topological phase transitions are a new class of transitions beyond Landau’s symmetry breaking theory, characterized by quantized, non-local topological invariants Hasan and Kane (2010); Qi and Zhang (2011). Recently, the non-equilibrium dynamics of topological phase transitions have received widespread attention both experimentally and theoretically Fläschner et al. (2018); D’Alessio and Rigol (2015); Hauke et al. (2014); Ye and Li (2020); Sun et al. (2018); Nie and Zheng (2023b), particularly concerning the critical behavior of quenching across topological phase transitions Liou and Yang (2018); Ulčakar et al. (2019, 2020); Sun et al. (2022); Kuo et al. (2021); Deng et al. (2024).
@@
Relating conventional continuous phase transitions and nonequilibrium dynamics, the Kibble-Zurek (KZ) mechanism was originally proposed by Kibble in his study of the formation of topological defects in the early universeKibble (1976), and it has been applied by Zurek to study phase transitions in condensed matter systems Zurek (1985); Zurek et al. (2005). It elucidates the universal scaling of a system when it is quenched across a continuous phase transition in a finite time $\tau$. In equilibrium, the theory of critical phenomena states that as the control parameters $V$ are tuned closer to their critical values $V_{\mathrm{c}}$, the correlation length $\xi$ and relaxation time $t_{\mathrm{r}}$ tend to diverge as $t_{\mathrm{r}}\sim|V-V_{\mathrm{c}}|^{-z\nu}$ and $\xi\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $z$ is the dynamic exponent and $\nu$ is the correlation length critical exponent. As the system slowly quenches across the transition, the relaxation time diverges as it approaches the critical point, and the system’s evolution ceases to be adiabatic. The time interval during which the system no longer evolves adiabatically is characterized as the freeze-out time $t_{\mathrm{f}}$. The freeze-out time and the unfrozen correlation length satisfy $t_{\mathrm{f}}\sim\tau^{z\nu/(1+z\nu)}$ and $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}$. The KZ behavior has been observed in diverse experimental systems, such as cosmic microwave background Bevis et al. (2008), superfluids Chesler et al. (2015); Mathey and Polkovnikov (2010), superconductor Sonner et al. (2015); Garaud and Babaev (2014); Monaco et al. (2006), liquid crystals Chuang et al. (1991), and ultracold atoms Donner et al. (2007); Weiler et al. (2008); Ko et al. (2019); Chen et al. (2011); Lamporesi et al. (2013); Navon et al. (2015); Clark et al. (2016); Keesling et al. (2019); Yi et al. (2020); Huang et al. (2021); Liang et al. (2022).
@@
Although topological transitions can not be described by Landau’s symmetry breaking theory, the KZ-like behavior has already been introduced to topological systems. Some studies have manifested that the total density of excitations shows a KZ power-law scaling Liou and Yang (2018); Ulčakar et al. (2019); Bermudez et al. (2009, 2010); Sim et al. (2022, 2023). However, this simple generalization does not reveal the underlying fundamental topological physics and shows incapability in extracting correlation lengths. Later, it is afterwards proposed that effective correlation length can be extracted from the Berry curvature profile in momentum space, and it shows power-law divergence at the transition Chen et al. (2017). Yet, this momentum-space description is not quite intuitive in contrast with the conventional KZ mechanism. Recently, significant breakthroughs have been made in the real-space analysis. It has been found that the local Chern marker (LCM) will show domain-like configurations when the system is disordered Ulčakar et al. (2020). As long as the disorder is weak, it is shown that the disorder-averaged normalized autocorrelation function of LCM can be regarded as an intrinsic property of the original clean topological system. Thus, the correlation length can be determined via this purely real-space method. During the quench, the domain-like configurations are the analogy to conventional topological defects and the correlation length shows KZ behavior. This method is first proposed in two-dimensional Qi-Wu-Zhang (QWZ) insulators Ulčakar et al. (2020) and then extended to one-dimensional Su-Schrieffer-Heeger (SSH) models Sun et al. (2022) by changing LCM to local winding number. However, since these two models have similar linear dispersion during gap closing, they have yielded the same critical exponents, $\nu\simeq 1,z\simeq 1$. Therefore, we naturally ask whether different critical exponents can be obtained.
@@
In this work, we systematically investigate the KZ behavior in the topological transition of a 2D checkerboard lattice model with a quadratic band crossing. For every scaling studied, dual perspectives are adopted: momentum-space Berry curvature in the clean system, and the real-space LCM in the disordered system. In the former, we mainly focus on the radial distribution of Berry curvature. In the latter, we are concerned with the domain-like configurations of the LCM and extract the real-space correlation lengths.
@@
Figure 1: (a). Checkerboard lattice model. Red (Blue) points denote the A (B) sublattices. NN hopping strength along the black arrows direction (opposite direction) is $J-\mathrm{i}V$ ($J+\mathrm{i}V$), which breaks the time-reversal symmetry as long as $V\neq 0$. Red (Blue) solid lines represent the NNN hopping strength $J_{1}^{\prime}(J_{2}^{\prime})$. (b). Topological phase diagram. Chern number $|C|=1$ acquires an opposite sign to $V$. At the transition, it is a time-reversal invariant semimetal. (c). The band structure of our model is in different regimes. The colors denote the polarization $\langle\sigma_{z}\rangle$. The gap closes with a quadratic dispersion at the transition $V=0$. We can see the band-inversion surface (BIS, shown as the white contours) is not a circle around symmetric momentum as in the QWZ model Zhang et al. (2018); Sun et al. (2018), which is a fundamental effect induced by the quadratic band crossing.
@@
We find that the length scale, in both perspectives, diverges near the critical point of the topological phase transition in equilibrium, satisfying $\xi\sim|V-V_{\mathrm{c}}|^{-1/2}$. By slowly quenching through the topological phase transition over a quench time $\tau$, we observe that the freeze-out time $t_{\mathrm{f}}$, defined by the delay time of the relaxation of Hall response in the clean system or the divergence of correlation length in the disordered system, satisfies $t_{\mathrm{f}}\sim\tau^{1/2}$. The corresponding length scale at the freeze-out time $t_{\mathrm{f}}$ satisfies $\xi(t_{\mathrm{f}})\sim\tau^{1/4}$, where the dynamic exponent $z\simeq 2$ and correlation length critical exponent $\nu\simeq 1/2$ differs from the results of linear band dispersion. Additionally, we extend the analysis to higher-order dispersions beyond the quadratic band crossing and find that the KZ power-law relations still hold. The corresponding critical exponents are related to the band-crossing dispersion.
@@ II Model
Two-dimensional (2D) checkerboard lattice Sun et al. (2009, 2011) is an example of a topological model with a quadratic band crossing. We consider a simplified checkerboard lattice model consisting of two sublattices (A and B) with real and isotropic nearest-neighbor (NN) hoppings ($J$), two kinds of next-nearest-neighbor (NNN) hoppings ($J_{1}^{\prime},J_{2}^{\prime}$), and complex NN hopping ($\mathrm{i}V$) leading to a topological nontrivial phase (see Fig.1(a)). The Hamiltonian can be written as:
$\begin{split}\hat{H}&=\sum_{\mathbf{r}}|x,y\rangle\langle x,y|\otimes\begin{pmatrix}0&J-\mathrm{i}V\\ J+\mathrm{i}V&0\end{pmatrix}\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x+1,y|\otimes\begin{pmatrix}J_{2}^{\prime}&0\\ J-\mathrm{i}V&J_{1}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y\rangle\langle x,y+1|\otimes\begin{pmatrix}J_{1}^{\prime}&J+\mathrm{i}V\\ 0&J_{2}^{\prime}\end{pmatrix}+\text{H.c.}\right]\\ &+\sum_{\mathbf{r}}\left[|x,y+1\rangle\langle x+1,y|\otimes\begin{pmatrix}0&0\\ J+\mathrm{i}V&0\end{pmatrix}+\text{H.c.}\right],\\ \end{split}$ (1)
where $\mathbf{r}=(x,y)$ denote the unit cell position (a unit cell is shown as a dashed ellipse in Fig.1(a)), and the $2\times 2$ matrices represent the sublattice tunneling related to the unit cell hopping.
After a Fourier transformation, we can express the Hamiltonian in momentum space ^$\hat{\mathcal{H}}_{\mathbf{k}}$ with momentum $\mathbf{k}=(k_{x},k_{y})$ as
^$\begin{split}\hat{\mathcal{H}}_{\mathbf{k}}=&\left[-\left(J_{1}^{\prime}+J_{2}^{\prime}\right)\left(\cos k_{x}+\cos k_{y}\right)\right]\hat{I}\\ &-4J\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\right)\hat{\sigma}_{x}-2V\left(\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\right)\hat{\sigma}_{y}\\ &-\left(J_{1}^{\prime}-J_{2}^{\prime}\right)\left(\cos k_{x}-\cos k_{y}\right)\hat{\sigma}_{z},\end{split}$ (2)
where $\hat{I}$ and $\hat{\sigma}_{x,y,z}$ are the identity and Pauli matrices, respectively. We set $J=1$, $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}$ and periodic boundary condition for the entire text afterwards. It is easy to read from Eq.2 that the system is an insulator without time-reversal symmetry as long as $V\neq 0$. Actually, the system has two topologically non-trivial phases, whose Chern number $C$ is determined by the sign of complex hopping $V$, as shown in Fig.1(b). Additionally, the energy band structures in different regimes are depicted in Fig.1(c). We can see that the energy gap closes at $\mathbf{G}=(k_{x},k_{y})=(\pi,\pi)$ when $V=V_{\mathrm{c}}=0$, exhibiting a quadratic band crossing with Chern number changing by $2$. This is totally different from the QWZ model. There, the gap closes at a Dirac cone, and every Dirac cone changes the Chern number with $1$. Besides, the band-inversion surface (BIS, defined as the $\langle\sigma_{z}\rangle=0$ contour) in the checkerboard lattice model forms a crisscross, while in the QWZ model, it is a circle around the band crossing point Zhang et al. (2018); Sun et al. (2018).
@@ III Ground state
For conventional second-order phase transitions driven by spontaneous symmetry breaking, the correlation lengths exhibit a universal scaling behavior as the systems approach the critical point. When it comes to topological transitions, such as the QWZ model and the SSH model, a scaling of $\nu\simeq 1$ has been found. However, can this exponent be different in other models with different band crossings?
@@ III.1 Perspective I: Clean system
Figure 2: (a). Correlation length in the ground state could be determined by the FWHM of the radial distribution of Berry curvature, here $J^{\prime}=1$. The closer to the transition, the narrower the peak. (b). This correlation length diverges in a power $\nu\simeq 1/2$ as the system approaches the phase transition.
@@
Some studies have provided correlation functions for topological phases, defined as the Fourier transform of the Berry curvature Chen et al. (2017); Chen and Schnyder (2019); Molignini et al. (2023), which measures the itinerant-circulation correlation between Wannier states Thonhauser et al. (2005); Marzari et al. (2012). Alternatively, a more direct definition of correlation length can also be extracted from the width of the profile of the Berry curvature, which will be explained in the following. Similar procedures have been used in studying the critical behavior in the QWZ insulator and the SSH model Ulčakar et al. (2020); Sun et al. (2022).
@@
The valence band Berry curvature $F(\mathbf{k})$ can be expressed via Berry connection $A_{k_{\alpha}}=\left\langle u_{\mathbf{k}}\left|\mathrm{i}\partial_{k_{\alpha}}\right|u_{\mathbf{k}}\right\rangle$ ($\alpha=x,y$) as
$F(\mathbf{k})=\partial_{k_{x}}A_{k_{y}}-\partial_{k_{y}}A_{k_{x}},$ (3)
@@
where $\left|u_{\mathbf{k}}\right\rangle$ is the valence band Bloch wavefunction at momentum $\mathbf{k}$. The Chern number mentioned before can be calculated by $C=\frac{1}{2\pi}\int_{\rm{FBZ}}F(\mathbf{k})\rm{d}^{2}\mathbf{k}$, where ”FBZ” represents the first Brillouin zone.
@@
Then, by integrating the Berry curvature angularly, $f(k)=\frac{1}{k}\int_{|\mathbf{k}^{\prime}-\mathbf{G}|=k}F(\mathbf{k}^{\prime})|\text{d}\mathbf{k}^{\prime}|$, we obtain its radial distribution $f(k)$. Fig.2(a) illustrates $f(k)$ with fixed $J^{\prime}=1$ and six different values of complex NN hoppings strength $V$. It is evident that the radial distribution of the Berry curvature exhibits a peak. Initially, this peak width decreases with the increasing $V$, reaching its minimum at the topological phase transition point $V_{\mathrm{c}}=0$, and subsequently increases. This reflects the critical behavior near the topological phase transition point.
@@
To describe this critical behavior quantitatively, we define the correlation length $\xi_{k}$ by the inverse of the full width at half maximum (FWHM) value of the radial distribution of Berry curvature $f(k)$. A similar approach has been utilized in the QWZ insulators and the SSH model Ulčakar et al. (2020); Sun et al. (2022). As shown in Fig.2(b), we observe that as the system approaches the critical point $V_{\mathrm{c}}=0$, the correlation length diverges. We plot the relationship between the correlation length and $|V-V_{\mathrm{c}}|$ in log-log coordinates under three sets of parameters, namely $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=0.5$, $1$, $1.5$. We find that the correlation length satisfies $\xi_{k}\sim|V-V_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/2$. This critical exponent value is distinct from the case of linear dispersion, such as in the QWZ insulators where $\nu\simeq 1$.
@@
Figure 3: (a). The real-space distribution of the LCM in a disordered sample with size $N\times N=50\times 50$. The area of the fragment domain increases to the order of the whole system sizes at the transition. (b). Divergence of real-space correlation length $\xi_{r}$ at transition with different system sizes in a log-log scale. Each data is an average of $20$ disorder samples. For a certain system size $N$, the correlation length shows good power law in $64<\xi<N/$ 6 regime (emphasized by the markers). The red line represent a $\left|V-V_{\mathrm{c}}\right|^{-1/2}$ scaling for reference. (c). The fitted power $\nu$ approaches $1/2$ as the system size increases. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
III.2 Perspective II: Disordered system